1/9 of 6/5 is 1/9 of 6/5 is 1/9 of 6/5 is 1/9 of 6/5 is

 1/9 of 6/5 is                                       
 1/9 of 6/5 is                                       
 1/9 of 6/5 is                                       
 1/9 of 6/5 is                                       

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

1/9 of 6/5 is 2/15.

= (1/9) × (6/5)
= (1/3) × (2/5)
= 2/15


Thanks